Please help me solve this problem: Multiply and divide. 3/5 times 5 1/2
Answer by Cromlix(4381)

Hi there,
3/5 x 5 1/2
5 1/2 = 11/2
3/5 x 11/2
= 33/10
= 3 3/10
Hope this helps :-)